OpenStudy (anonymous):

Abel throws a dart that hits the square shown below: A square is drawn. A circle of radius 4 units touches the sides of the square. What is the probability that the dart does not hit a point in the circle? 16% 21.5% 64% 78.5%

OpenStudy (anonymous):

@Michele_Laino @mathrulezz

OpenStudy (irishboy123):

this problem assumes that he dart has the same odds of hitting anywhere inside the square. on that basis, just divide the area of circle by area of square.

OpenStudy (anonymous):

Oh and how do you do that? @IrishBoy123

OpenStudy (irishboy123):

πr^2/(r+r)^2

OpenStudy (anonymous):

area of square: 8x8 = 64

OpenStudy (anonymous):

so then 4 x 2/8 x 2?

OpenStudy (anonymous):

8/16 or 1/2 then, right?

OpenStudy (anonymous):

@IrishBoy123 last question..... is it A?

OpenStudy (irishboy123):

πr^2/(r+r)^2 = π/4 = 0.785, thus 78.5% of the area of square is covered by circle. BUT you wish to calc the odds dart DOES NOT hit point in circle.

OpenStudy (anonymous):

oh, thx :) medal and fan:)

OpenStudy (anonymous):

there u go :)

OpenStudy (anonymous):

bye ;-)

OpenStudy (irishboy123):

cool